OpenStudy (xochital):

What is the value of the x variable in the solution to the following system of equations? 2x + 3y = 4 x - 2y = -5 2 -2 1 -1

OpenStudy (anonymous):

because you get a set of answer, just try each of them in for the x and see which has the same solution. if x=2 then y=0 and y=-7/5 X if x=-2 then y=8/3 and y=3/2 X if x=1 then y=3/2 and y=3 X if x=-1 then y=2 and y=2 O the x solution is -1 and the y solution is 2.

OpenStudy (anonymous):

the alternate method of solving this would be to isolate one variable and insert it into the other eqation. x=2y-5 can be assumed given x-2y=-5 2(2y-5) + 3y =4 <--- this can be solved for the Y value. w/e that is can be inserted into either equation for the X value.

OpenStudy (xochital):

thank you!! :) can you help me with another?

OpenStudy (anonymous):

sure

OpenStudy (xochital):

For the following system, if you isolated x in the second equation to use the Substitution Method, what expression would you substitute into the first equation? 3x + y = 8 -x - 2y = -10 -2y + 10 2y + 10 2y - 10 -2y - 10

OpenStudy (anonymous):

to isolate X in the second equation would give you: \[x=-2y+10\] that is substituted into the first to replace X. -2y+10 is you answer

OpenStudy (xochital):

thank you so much! :D

OpenStudy (anonymous):

x = -1 & y = 2